Show that the value of the integral

∫ [(x^2+3y)dx + (5x-3y^2)dy]

C

where C is the ellipse x^2/a^2+y^2/b^2=1, is twice the area enclosed by C.


1
Expert's answer
2021-06-01T06:13:37-0400

The above equation is -


"\int_{c}[(x^{2}+3y)dx+(5x-3y^{2})dy]" ,


where C is the ellipse "\dfrac{x^{2}}{a^{2}}+\dfrac{y^{2}}{b^{2}}=1"


Now comparing this equation with green's theorem , we can only solve this question by green's theorem.



Now statement of green's theorem is that ,


"\oint_{c}(Mdx+Ndy)=\iint_{R}(\dfrac{\partial N}{\partial x}-\dfrac{\partial M}{\partial y})dxdy" ...............A)



now this is the green's theorem now we will compare the given equation with the given equation , we will get -


"M=x^{2}+3y"


"N=5x-3y^{2}"


"\dfrac{\partial M}{\partial y}" "=3"


"\dfrac{\partial N}{\partial x}=5"



Now putting the value in equation A) , we get our right hand side as -


"=2\iint_{R}dxdy"


"where" "\iint_{R}dxdy" "=area\ of\ ellipse"


"=2(" area of ellipse)


"=" area of given ellipse ="{\pi}ab"


"=" so , our required answer is "=2{\pi}ab"



which is the twice the area enclosed by C.